# Probing the fission of 220Ra⋆ at E⋆ 31.8–45.4 MeV

### v1

@@ 4.2 Isotopic yield distribution
In heavy composite systems, the emission of neutrons competes with the process of fission, particularly at moderate excitation energies. However, the emission of charged particles faces a significant challenge due to the presence of the Coulomb barrier. As a result, neutron emission becomes a competitive process against fission. The emission of nucleons from the primary fission fragments and/or the fission of precursors of the fission-decay chains can lead to isotopic and isobaric yield distributions of the fission residues. The fact that neutron emission is more probable than proton emission results in the observation of different isotopes in most cases. An isotopic chain can arise from the sequential evaporation of neutron(s) from fission fragments, yielding a fission product with a fixed atomic number and mass number .
Four isotopes of antimony (115,116m,117,118mSb) and three isotopes of indium (109,110,111In) have been identified at E⋆ = 45.4 and 39.6 MeV, and at E⋆ = 45.4 MeV, respectively. The experimentally measured yields of Sb and In isotopes have been analyzed as per the prescription presented in Ref. [46] to achieve isotopic yield distribution and to obtain the charge distribution parameters. In the case of Sb, the two isomers are high spin relative to the ground spin; the cross-section of the high spin isomer has been measured, or the high spin isomer completely decayed to the low spin isomer through IT (Isomeric Transition). The measured yields of the Sb isotopes have been approximated as the total cross-section of the respective isotopes. The parameters of isotopic yield distribution, e.g., most probable mass p and the width parameter , for antimony and indium isotopes at E⋆ = 45.4 and 39.6 MeV were obtained by fitting a Gaussian function to their respective yields, and are shown in Figs. 6(a)- 6(c), respectively. The values of most probable mass p for antimony isotopes at excitation energy E⋆ = 45.4 and 39.6 MeV estimated to be 116.027 0.014 and 116.029 0.008 MeV, and the width parameters for isotopic yield distribution are estimated to be 1.711 0.012 and 1.628 0.005, respectively. For indium isotopes, the values of p and at E⋆ = 45.4 MeV is found to be 110.178 0.005 and 1.113 0.008. The uncertainties in these estimated parameters are the fitting errors. The values of p and 2 of Sb and In isotopes at E⋆ = 45.4 and 39.6 MeV are reported in Table 4.
| E⋆ | Isotope | Most Probable mass | Isotopic width |
| (MeV) | p | 2 | |
| 45.4 | Sb | 116.027 0.014 | 3.422 |
| 39.6 | Sb | 116.029 0.008 | 3.256 |
| 45.4 | In | 110.178 0.005 | 2.226 |
| System | E⋆ (MeV) | Isotope | Ref. | |
@@
| Target | Projectile | Isotope | ||
@@
The variance parameter () for Sb isotopes at excitation energy E⋆ = 45.4 and 39.6 MeV is found to be 2.93 and 2.65, and for In isotopes is found to be 1.24 at E⋆ = 45.4 MeV, respectively. The value of for several other systems at comparable excitation energies is shown in Table 5 along with the present system. As can be noticed from this table, the value of for the present system agrees reasonably well with those reported in the literature. For better insights into the spread of isotopes in different systems (given in Table 5), the value of variance parameter () of different isotopes and the mass-asymmetry parameter () for various projectile-target combinations are plotted in Fig. 7. As shown in this figure, the value of variance for these isotopes falls within a range of width 4.13. The average value of the variance parameter is found to be 3.86, suggesting the involvement of fusion-fission dynamics in the population of Sb and In isotopes. It is in line with the observations reported in the literature. The data presented in Table 6 suggests the higher value of variance parameter () for less mass-asymmetric systems for the same target.
@@ 4.3 Isobaric yield distribution
Charge distribution describes how the nuclear charge is distributed between the two complimentary fission fragments for a given projectile-target combination undergoing fission. To obtain such distribution, it is essential to determine the most probable atomic number (p) for the fission product with the highest yield among all the products of a given mass chain and the fractional yield (FY). The value of FY for an isotope is calculated by dividing the yield of the isotope by the total yield of the mass chain to which it belongs. The most probable charge p for a given isotopic distribution can be calculated using a formula based on the Unchanged Charge Distribution (UCD) hypothesis [52] as,
where and represent the charge and mass of the fission fragment, respectively, while p signifies the most probable mass. The value of FY, obtained by normalizing experimentally determined yields, of Sb at E⋆ = 45.4 and 39.6 MeV, and In at E⋆ = 45.4 MeV, are plotted as a function of charge corrected isotopic fragments ( - p) in Figs. 8(a)- 8(c), respectively. The charge dispersion parameter obtained from the Gaussian fit to the data for Sb is estimated to be 0.769 and 0.714 at E⋆ = 45.4 and 39.6 MeV, respectively. For In isotopes, the value of is estimated to be 0.430 at E⋆ = 45.4 MeV.
Additionally, the charge dispersion parameter () has been calculated by converting the variance parameter () of isotopic yield distribution into () using the formulation given below,
The values of obtained for different isotopes as a function of energy are reported in Table 7. As can be seen from this table, the values of obtained from the analysis of charge distribution data presented in Figs. 8(a)- 8(c) and the one calculated from the above formulations, are found to be in good agreement, indicating the self-consistency of the approaches used in the present analysis.
| E⋆ | Isotope | z | z |
| (MeV) | From Figs. 8(a)- 8(c) | From Eq.(2) | |
@@
Additionally, the charge dispersion parameter () has been calculated by converting the variance parameter () of isotopic yield distribution into () using the formulation given below,
The values of obtained for different isotopes as a function of energy are reported in Table 7. As can be seen from this table, the values of obtained from the analysis of charge distribution data presented in Figs. 8(a)- 8(c) and the one calculated from the above formulations, are found to be in good agreement, indicating the self-consistency of the approaches used in the present analysis.
@@ 5 Summary and conclusions
The production cross-sections of various fission fragments populated in the 12C+208Pb system have been measured at E⋆ = 45.4, 39.6, and 31.8 MeV. The mass distribution of fission fragments has been studied at different excitation energies to investigate the dispersion of fission fragments. The presence of asymmetric fission suggests the role of nuclear shell effects in the studied excitation energies. The mass distribution data have been analyzed to obtain the isotopic and isobaric yield distributions of antimony (Sb) and indium (In) isotopes. It has been found that a single Gaussian function effectively explains the isotopic yield distributions for Sb and In isotopes. The derived mass and charge dispersion parameters agree reasonably well with the experimental values reported in the literature for other fissioning systems. The values of variance for many such distributions in literature and obtained in the present study for Sb and In fall within a narrow range of width 4.13. The data analysis further indicates that the entrance-channel mass asymmetry significantly affects the variance of isotopic yield distribution. The isobaric distribution has been obtained for Sb and In isotopes from the analysis of their experimental yields to get the charge dispersion parameter . The value of for Sb was found to be 0.769 and 0.714 at E⋆ = 45.4 and 39.6 MeV, respectively. For In isotopes, was found to be 0.430 at E⋆ = 45.4 MeV. These values of show a good agreement with the values of calculated by substituting the obtained isotopic distribution parameter in theoretical expression. The agreement between these values shows the self-consistency of the present analysis and the experimental data. In the present study, medically important 99mTc and 111In isotopes with decent production probability have been observed. As such, the fission of excited 220Ra nuclei may provide an alternative path to producing 99mTc and 111In isotopes for application in radio medicine. For better insights into the fission dynamics of 220Ra, an online experiment to measure the neutron multiplicity is in order.

### v3

@@ 3.1 Isotopic yield distribution
In a heavy composite system, the emission of particles (charged and/or uncharged) competes with the process of fission, particularly at moderate excitation energies. Therefore, the emission of particles from the fission fragments may give rise to the isotopic and isobaric distributions of the fission residues. However, the emission of charged particles faces a significant challenge due to the presence of the Coulomb barrier. The fact that neutron emission is more probable than proton emission, results in the observation of isotopic yield distribution in most cases. The study of isotopic yield distribution insights into the major characteristics of low energy fission [37].
In the present work, four isotopes of antimony (115,116,117,118mSb) and three isotopes of indium (109,110,111In) have been identified at E⋆ = 45.4 and 39.6 MeV, and at E⋆ = 45.4 MeV, respectively. The experimentally measured yields of Sb and In isotopes have been analyzed as per the prescription presented in Ref. [38] to achieve isotopic yield distribution and to obtain the charge distribution parameters. The parameters of isotopic yield distribution, e.g., most probable mass p and width parameter , for antimony and indium isotopes at E⋆ = 45.4 and 39.6 MeV were obtained by fitting a Gaussian function to their respective yields as shown in Figs. 4(a)- 4(c), respectively. The values of most probable mass p for antimony isotopes at excitation energy E⋆ = 45.4 and 39.6 MeV estimated to be 116.027 0.014 and 116.029 0.008 MeV, and the width parameters for isotopic yield distribution are estimated to be 1.711 0.012 and 1.628 0.005, respectively. For indium isotopes, the values of p and at E⋆ = 45.4 MeV is found to be 110.178 0.005 and 1.113 0.008. The uncertainties in these estimated parameters are the fitting errors.
| System | E⋆ (MeV) | Element | Ref. | |
@@
| Target | Projectile | Element | ||
@@
The variance parameter () for Sb isotopes at excitation energy E⋆ = 45.4 and 39.6 MeV is found to be 2.93 and 2.65, and for In isotopes is found to be 1.24 at E⋆ = 45.4 MeV, respectively. The value of for several other systems at comparable excitation energies is shown in Table 3 along with the present system. As can be noticed from this table, the value of for the present system agrees reasonably well with those reported in the literature and give confidence to the present observations. For better insights into the spread of isotopes in different systems (given in Table 3), the value of variance parameter () of different isotopes and the mass-asymmetry parameter () for various projectile-target combinations are plotted in Fig. 5. As shown in this figure, the value of variance for these isotopes falls within a range of width 4.13. The average value of the variance parameter is found to be 3.86, suggesting the involvement of fusion-fission dynamics in the population of Sb and In isotopes. It is in line with the observations reported in the literature. The data presented in Table 4 suggests a higher value of variance parameter () for less mass-asymmetric systems for the same target.
@@ 3.2 Isobaric yield distribution
The isobaric charge distribution is a crucial post-fission observable. As demonstrated by Gubbi et al. [20], to deduce the total chain yield of a given fission residue of mass A, it is essential to have the information on the isobaric charge dispersion parameter () and the most probable charge (p) for the fission product with the highest yield among all the products of a given mass chain . The p for the antimony and indium isotopes are calculated using the relation:
where and represent the charge and mass of the fission-like fragments, respectively, while p signifies the most probable mass. In this study, an effort has been made to analyze the isobaric charge distribution in terms of fractional independent yield (FIY). The fractional independent yields for isotopes Sb and In were obtained by dividing the independent yields by their corresponding charge yields. The distribution of fractional chain yield vs charge corrected isotopic fragments ( - p) is shown in Figs. 6(a)- 6(c), respectively. The charge dispersion parameter () obtained from the Gaussian fit to the data for Sb is estimated to be 0.769 and 0.714 at E⋆ = 45.4 and 39.6 MeV, respectively. For In isotopes, the value of is estimated to be 0.430 at E⋆ = 45.4 MeV.
| E⋆ | Element | Z | Z |
| (MeV) | From Figs. 6(a)- 6(c) | From Eq.(2) | |
@@
Additionally, the charge dispersion parameter () has been calculated by converting the width parameter () of isotopic yield distribution into using the formulation given below,
The values of obtained for different isotopes as a function of energy are reported in Table 5. As can be seen from this table, the values of obtained from the analysis of charge distribution data presented in Figs. 6(a)- 6(c) and the one calculated from the above formulations, are found to be in good agreement, indicating the self-consistency of the approaches used in the present analysis.
3.3 Mass distribution of fission-like fragments
The mass distribution of fission fragments is a crucial post-fission observable directly linked to the collective dynamics of the fission [43]. The activities produced in the target-catcher assembly were used to determine the mass distribution. To generate mass distributions, the experimentally determined production cross-sections of fission-like fragments produced in the 12C + 208Pb system are plotted as a function of mass number (A) in Figs. 7(a)- 7(b) at E* = 45.4 and 39.6 MeV respectively. The upward arrows indicate that only the metastable states have been measured, and the total production cross-sections of these fission-like fragments are expected to increase. From these figures, it can be seen that the observed mass distribution is Gaussian-like and fitted with a Gaussian function suggesting the formation of these fission-like fragments from the compound nuclear process. The solid lines represent the Gaussian fit through data points. The present observation is in line with the observations reported by Pokrovsky et al. [31]. The centroid (Mp) and width () of the mass distribution curve at E* = 45.4 MeV are found to be 114.29 0.50 and 15.38 0.82 respectively. At E* = 39.6 MeV, the value of Mp and are found to be 113.56 0.32 and 12.99 0.56 respectively. The quoted uncertainties are the fitting errors. The value of full width at half maximum (FWHM) at E* = 45.4 and 39.6 MeV is found to be 36.22 and 30.61 respectively. The FWHM of the distributions is comparable with those observed in 20Ne + 209Bi [44], 20Ne + 208Pb [21], 16O + 208Pb [30]. The fission dynamics are significantly impacted by the predominant direction of mass flow within the dinuclear system. As reported in the literature [45, 46, 47], if the mass asymmetry () is greater than the Businaro-Gallone critical mass asymmetry (BG) [48], mass flows from the projectile to the target, leading to the formation of the compound nucleus which may eventually decay via the process of fission. If the value of , the mass flow occurs from the target to the projectile, and a dinuclear system is formed, decaying before equilibrating in all degrees of freedom. For 12C + 208Pb system, the critical mass asymmetry () is 0.86. In contrast, the mass asymmetry () is 0.89, i.e., , resulting in the occurrence of fission via compound nucleus formation path.
In order to study the change in mass variance () with excitation energy, the value of is plotted as a function of excitation energy in Fig. 8. The value of increases with excitation energy, indicating a large spread in the fission fragment masses at higher excitation energies. The observed variation in the value of with excitation energy above the barrier follows the same trend as reported by Ghosh et al. [17]. It is important to mention that this study only examined the system at two energies above the barrier. Consequently, additional research is required to explore how the value of changes with excitation energy at and below the barrier energies for a more comprehensive understanding.
@@ 4 Summary and conclusions
The production cross-sections of various fission-like fragments populated in the 12C + 208Pb system have been measured at E⋆ = 45.4 and 39.6 MeV. The isotopic and isobaric yield distributions of two fission-like fragments antimony (Sb) and indium (In) have been obtained. It has been found that a single Gaussian function effectively explains the isotopic yield distributions for Sb and In isotopes. The derived mass and charge dispersion parameters agree reasonably well with the experimental values reported in the literature for other fissioning systems. The values of variance for many such distributions in literature and obtained in the present study for Sb and In fall within a narrow range of width 4.13. The data analysis further indicates that the entrance-channel mass asymmetry significantly affects the variance of isotopic yield distribution. The isobaric distribution has been obtained for Sb and In isotopes from the analysis of their experimental yields to get the charge dispersion parameter . The value of for Sb was found to be 0.769 and 0.714 at E⋆ = 45.4 and 39.6 MeV, respectively. For In isotopes, was found to be 0.430 at E⋆ = 45.4 MeV. These values of show a good agreement with the values of calculated by substituting the obtained isotopic distribution parameter in theoretical expression. The agreement between these values shows the self-consistency of the present analysis and the experimental data. The mass distribution of fission-like fragments has been studied at different excitation energies to investigate the dispersion of fission-like fragments. The mass distributions at two studied energies are found to be symmetric and fitted with a Gaussian function, indicating the population of these fission-like fragments via deexcitation of the compound nucleus. It has been observed that the mass variance increases with increases in excitation energy. In the present study, medically important 99mTc and 111In isotopes with decent production cross-section have been observed. As such, the fission of excited 220Ra nuclei may provide an alternative path to producing 99mTc and 111In isotopes for application in radio medicine. For better insights into the fission dynamics of 220Ra, an online experiment to measure the neutron multiplicity is in order.